OBERSEMINAR ÜBER GEOMETRISCHE STOCHASTISCHE ANALYSIS

K.-T. STURM, A. THALMAIER WS 1999/00

References

[1] S. Aida, Sobolev spaces over loop groups. J. Funct. Anal. 127 (1995), 155–172.

[2] S. Aida, Logarithmic Sobolev inequalities on loop spaces over compact Riemannian manifolds. Stochastic analysis and applications (Powys, 1995), 1–19, World Sci. Publishing, River Edge, NJ, 1996.

[3] S. Aida, Uniform positivity improving property, Sobolev inequalities, and spectral gaps. J. Funct. Anal. 158 (1998), 152–185.

[4] S. Aida, Differential calculus on path and loop spaces. II. Irreducibility of Dirichlet forms on loop spaces. Bull. Sci. Math. 122 (1998), 635–666.

[5] S. Aida and K. D. Elworthy, Differential calculus on path and loop spaces. I. Logarithmic Sobolev inequalities on path spaces. C. R. Acad. Sci. Paris Sér. I Math. 321 (1995), 97–102.

[6] H. Airault and P. Malliavin, Integration on loop groups. II. Heat equation for the Wiener measure. J. Funct. Anal. 104 (1992), 71–109.

[7] S. Albeverio, R. Léandre and M. Röckner, Construction of a rotational invariant diffusion on the free loop space. C. R. Acad. Sci. Paris Sér. I Math. 316 (1993), 287–292.

[8] M. Capitaine, E. P. Hsu and M. Ledoux, Martingale representation and a simple proof of logarithmic Sobolev inequalities on path spaces. Electron. Comm. Probab. 2 (1997), 71-81.

[9] A.-B. Cruzeiro and P. Malliavin, Renormalized differential geometry on path space: structural equation, curvature. J. Funct. Anal. 139 (1996) 119–181.

[10] B. K. Driver, A Cameron-Martin type quasi-invariance theorem for Brownian motion on a compact Riemannian manifold. J. Funct. Anal. 110 (1992) 272–376.

[11] B. K. Driver, Loop space, curvature, and quasi-invariance. Centro Vito Volterra Lecture Notes, Universita Degli Studi Di Roma II, 1992.

[12] B. K. Driver, Cameron-Martin type quasi-invariance theorem for pinned Brownian motion on a compact Riemannian manifold. Trans. Amer. Math. Soc. 342 (1994), 375–395.

[13] B. K. Driver, Towards calculus and geometry on path spaces. Stochastic analysis (Ithaca, NY, 1993), 405–422, Proc. Sympos. Pure Math. 57, Amer. Math. Soc., Providence, RI, 1995.

[14] B. K. Driver, A Primer on Riemannian Geometry and Stochastic Analsysis on Path Spaces. Lecture notes of talks given at the ETH (Zürich, Switzerland), 1995.

[15] B. K. Driver, Integration by parts and quasi-invariance for heat kernel measures on loop groups. J. Funct. Anal. 149 (1997), 470–547.

[16] B. K. Driver, Integration by parts for heat kernel measures revisited. J. Math. Pures Appl. 76 (1997), 703–737.

[17] B. K. Driver and T. Lohrenz, Logarithmic Sobolev inequalities for pinned loop groups. J. Funct. Anal. 140 (1996), 381–448.

[18] B. K. Driver and M. Röckner, Construction of diffusions on path and loop spaces of compact Riemannian manifolds. C. R. Acad. Sci. Paris Sér. I Math. 315 (1992), 603–608.

[19] A. Eberle, Diffusions on path and loop spaces: existence, finite-dimensional approximation and Hölder continuity. Probab. Theory Related Fields 109 (1997), 77–99.

[20] K. D. Elworthy and X.-M. Li, A class of integration by parts formulae in stochastic analysis. I. Itô’s stochastic calculus and probability theory, 15–30, Springer, Tokyo, 1996.

[21] O. Enchev and D. W. Stroock, Towards a Riemannian geometry on the path space over a Riemannian manifold. J. Funct. Anal. 134 (1995) 392–416.

Date: October 2, 1999.

[22] O. Enchev and D. W. Stroock, Integration by parts for pinned Brownian motion. Math. Res. Lett. 2 (1995), 161–169.

[23] S. Fang, Stochastic anticipative calculus on the path space over a compact Riemannian manifold. J. Math. Pures Appl. 77 (1998), 249–282.

[24] S. Fang, Integration by parts formula and logarithmic Sobolev inequality on the path space over loop group. Ann. of Prob. 27 (1999), 664–683.

[25] S. Fang and J. Franchi, De Rham-Hodge-Kodaira operator on loop groups. J. Funct. Anal. 148 (1997), 91–407.

[26] S. Fang and P. Malliavin, Stochastic analysis on the path space of a Riemannian manifold. I: Markovian stochastic calculus. J. Funct. Anal. 118 (1993) 249–274.

[27] F.-Z. Gong and Z.-M. Ma, The log-Sobolev inequality on loop space over a compact Riemannian manifold. J. Funct. Anal. 157 (1998), 599–623.

[28] E. P. Hsu, Quasi-invariance of the Wiener measure on the path space over a compact Riemannian manifold. J. Funct. Anal. 134 (1995) 417–450.

[29] E. P. Hsu, Flows and quasi-invariance of the Wiener measure on path spaces. Stochastic analysis (Ithaca, NY, 1993), 265–279, Proc. Sympos. Pure Math. 57, Amer. Math. Soc., Providence, RI, 1995.

[30] E. P. Hsu, Inégalités de Sobolev logarithmiques sur un espace de chemins. C. R. Acad. Sci. Paris Sér. I Math. 320 (1995), 1009–1012.

[31] E. P. Hsu, Logarithmic Sobolev inequalities on path spaces. New trends in stochastic analysis (Charingworth, 1994), 168–181, World Sci. Publishing, River Edge, NJ, 1997.

[32] E. P. Hsu, Integration by parts in loop spaces. Math. Ann. 309 (1997), 331–339.

[33] E. P. Hsu, Logarithmic Sobolev inequalities on path spaces over Riemannian manifolds. Comm. Math. Phys. 189 (1997), 9–16.

[34] E. P. Hsu, Analysis on Path and Loop Spaces. In: Probability Theory and Applications, ed. by Elton P. Hsu and S. R. S. Varadhan. Amer. Math. Soc., Providence, RI, 1999.

[35] E. P. Hsu, Estimates of derivatives of the heat kernel on a compact Riemannian manifold. Proc. Amer. Math. Soc., to appear.

[36] T. Kazumi, Le processus d’Ornstein-Uhlenbeck sur l’espace des chemins riemanniens et le problème des martingales. J. Funct. Anal. 144 (1997), 20–45.

[37] M. Ledoux, Concentration of measure and logarithmic Sobolev inequalities. Manuscript, Uni-

versité Paul-Sabatier, Toulouse, 1998. http://www-sv.cict.fr/lsp/Ledoux/index.html

[38] T. Lyons and Z.-M. Qian, A class of vector fields on path spaces. J. Funct. Anal. 145 (1997) 205–223.

[39] P. Malliavin, Stochastic analysis. Grundlehren der Mathematischen Wissenschaften, 313. Springer-Verlag, Berlin, 1997.

[40] M.-P. Malliavin and P. Malliavin, Integration on loop groups. I. Quasi invariant measures. J. Funct. Anal. 93 (1990), 207–237.

[41] M.-P. Malliavin and P. Malliavin, Integration on loop group. III. Asymptotic Peter-Weyl orthogonality. J. Funct. Anal. 108 (1992), 13–46.

[42] P. Malliavin and D. W. Stroock, Short time behavior of the heat kernel and its logarithmic derivatives. J. Differential Geom. 44 (1996), 550–570.

[43] I. Shigekawa, Differential calculus on a based loop group. In: New Trends in Stochastic Analysis, Proceedings of the Taniguchi International Symposium, ed. by K. Elworthy, S. Kusuoka and I. Shigekawa, pp. 375-398. World Scientific, Singapore, 1997.

[44] D. W. Stroock, Some thoughts about Riemannian structures on path space. Gaz. Math. 68 (1996), 31–45.

[45] D. W. Stroock, An estimate on the Hessian of the heat kernel. Itô’s stochastic calculus and probability theory, 355–371, Springer, Tokyo, 1996.

[46] D. W. Stroock and J. Turetsky, Short time behavior of logarithmic derivatives of the heat kernel. Asian J. Math. 1 (1997), 17–33.